It is given that 


"z+2i=iz+\lambda,\quad\quad" "\frac{w}{z}=2+2i,\quad\quad" "\mathrm{Im}w=8,"

where z and w are complex numbers, and λ is a real constant. Which of the following is the value of \lambdaλ?



λ=4


λ=8


λ=3


λ=−3


λ=−4


1
Expert's answer
2021-10-19T17:19:02-0400

"z+2i=iz+\lambda\\\frac{w}{z}=2+2i\\z(1-i)=\lambda-2i\\z=\frac{\lambda-2i}{1-i}"

Rationalise the denominator

"z=\frac{(\lambda-2i)(1+i)}{(1)^2-(-1)^2}=\frac{\lambda+\lambda i-2i-2i^2}{2}"


"z=\frac{\lambda+2}{2}+\frac{i(\lambda-2)}{2}"


"z=\frac{\lambda+2}{2}+\frac{i(\lambda-2)}{2}"


Given, imaginary part of w=8)


"w=(2+2i)z"


"=(2+2i)\frac{(\lambda+2)+i(\lambda-2)}{2}"


"=(1+i)[(\lambda+2)+i(\lambda-2)]"


"=\lambda+2+i(\lambda-2)+i(\lambda+2)-1(\lambda-2)\\=\lambda+2-\lambda+2+i(\lambda-2+\lambda+2)\\=4+i(2\lambda)"


Since, ln(w)=8

"2\lambda=8\\\lambda=4"